Use an iterated integral to find the area of the region bounded by the graphs of the equations y = 27- xand y = x +7.)
Answer #1

$ y = x + 7 27 MARSEILLEUR 7=27-3? Intersecting of x +7= 27-22 x²+x-2020 (2-4)(x+5)=0 x = 4,5 PASSA I tis A 64 - Tilt Area ofAnswer question